### v3

# Regular triangle unions with maximal number of sides

[…]
2025
Regular triangle unions with maximal number of sides
[…]
Intro
Latvian, Lithuanian and (now extinct) Old Prussian constitute a Baltic branch of Indoeuropean (IDE) languages. During a Baltic Culture Festival in Northern Lithuania, August 2024, the author fixed his eyes on a scene decoration, where Latvian national symbol Auseklis [^1] From Latvian aust, Lithuanian aušti, ‘to dawn’, which has common IDE roots with English East. was shown. It forms an octagonal star (Fig. 1, left), which is just a standard octagram ($\{8,3\}$-gon in Schläfli notation). Though this cross is supposed to be drawn without lifting a stick from the ground (the reason behind this restriction is rooted, of course, in magic; same picture, left, where simple and double Lietuvēns crosses are shown), one clearly can distinguish $4$ triangles whose union forms a simple $16$-gon. With few more thoughts, it is immediate how one can increase edge count by slightly deforming all triangles. It was therefore natural to wonder what is the best possible result. Such is a subjective genesis of the problem. And magic surely works! The author could not anticipate that hypostases of both crosses will eventually emerge in this purely mathematical endeavour. Yet, the earliest to rise will be $\ldots$ the Star of David. But let us go to the basics first.
Figure 1: Left – a stamp dedicated to Latvian national Front (a movement that eventually led to liberation from Soviet occupation) anniversary featuring Auseklis – a Latvian national symbol. Right – a double and simple Lietuvēns cross.
Unions
[…]
Take $n$ $\triangle$. Its edges will henceforth be called legs, while edge will be reserved solely for a $\bigcup$ polygon. Any two of $\triangle$ intersect in at most $6$ points, and it is straightforward to contrive an example with $\sim 2n^{2}$ edges on the boundary of the $\bigcup$. However, in such construction many of the $\triangle$ appear to be very slim. If all $\triangle$ happen to have angles at least $\delta>0$ (so called $\delta$-fat"), the complement of the $\bigcup$ then has at most $O_{\delta}(n)$ connected components, and the boundary consists of $O_{\delta}(n\log\log n)$ edges fat . What happens, though, if the $\bigcup$ has no holes at all and its interior is connected? Let us formulate this precisely.
Problem.
Given an integer $n\geq 1$. Suppose, a simple polygon is a $\bigcup$ of $n$ $\triangle$. What is the maximal number of sides it can have? The following also applies.
  • 1)
    No further restrictions. This defines seq. $i(n)$.
  • 2)
    Vertices of $\triangle$ along the border of the $\bigcup$ are arranged in a cyclic way – seq. $e(n)$.
  • 3)
    Vertices of $\triangle$ are on the unit circle – seq. $c(n)$.
  • 4)
    Vertices are on the unit circle and are arranged in the cyclic way – seq. $r(n)$.
It may seem surprising that the question “how many edges exactly?" has not been asked before. Yet, there are reasons for that. The seq. $i(n)$ is the one where existing techniques and results apply most. Take a $\bigcup$ of $n$ $\triangle$ with $i(n)$ edges. Find a $\triangle$ which contributes least to the common boundary. Remove it. It is not difficult to see this reduces the edge count of the outer border by at most $\frac{i(n)}{n}+3$. Thus, as if
$i(n-1)\geq i(n)-\frac{i(n)}{n}-3.$
This would imply $i(n)=O(n\log n)$. Sure, such a naïve approach is not strictly valid: we do not know whether removal of 11 $\triangle$ leaves no holes in the $\bigcup$. But a certain aspect of the phenomenon is captured. In the Section ‘Lower Envelopes’, employing the theorem of Micha Sharir Ady Wiernik (deep and unexpected result for any newcomer to the topic), we will see that one cannot expect to do much better than that: $i(n)=\Omega(n\alpha(n))$, where $\alpha(n)$ is the functional inverse of the so-called Ackermann function. It tends to $\infty$ with an argument, but extremely slowly. Its value is $3$ in the interval $[5,16]$, and $4$ in the interval
$\Bigg[17,\underbrace{2^{2^{\ldots 2^{2}}}}_{65536\text{ copies}}\,\Bigg].$
It is impossible for a sane mind to fathom a place on the line of positive integers where this function surpasses $5$. There is a handful of other reasons why a plain answer is not to be expected.
Figure 2: I=mnmn+3I=mn-m-n+3 number of intersections of $n$-gon and $m$-gon for odd $m,n$ ($m=9$, $m=7$, $I=50$ here).
However, the situation with cyclically arranged vertices is slightly different. Such $\triangle$ $\bigcup$ will be called regular. The current paper, though mostly expository and derivative, introduces such objects for the first tiome and gives a new additional tool in dealing with $\bigcup$. Namely, triangulation shifts. Theoretically, an answer becomes within a reach. Sure, there exists problems in combinatorial geometry where an arm-stretch distance is still a considerable one.
For example, it is known that two simple polygons with odd numbers ($m$ and $n$) of vertices can intersect in at least mnnm+3mn-n-m+3 points (Fig. 2), and this is conjectured to be the peak value. Ackerman et al. ackerman proved that the maximum is indeed mnmn+Cmn-m-n+C. Accordingly, this is optimal, apart from the exact value of $C$ (still not rigorously known to be equal to $3$).
Henceforth, we will concentrate on $e(n)$, referred here as the Politrigon seq.. It now has a number A375986 in the OEIS oeis . What are its initial values? For $n=1$, there is nothing to prove. The $\bigcup$ of 22 $\triangle$, if a simple $M$-gon, can have any number of sides in the range $3\leq M\leq 10$, and also $M=12$. The latter is mostly familiar from a pleasingly symmetric hexagram known as the star of David. In order to prove that $e(3)=22$, case-by-case analysis can be performed. The example ($n=1,2,3$) is presented in Fig. 11 (three brown $\triangle$).
Figure 3: Regular $\bigcup$ of $4$ triangles with $33$ sides.
Figure 4: $10$ lines forming $25$ non-overlapping $\triangle$.
Figure 5: All 44 perfect $\bigcup$ of triangles. These are exactly the configurations satisfying (n+1)e(n)=(12n6)n(n+1)e(n)=(12n-6)n. Allegorically speaking, the line of succession to the Eye of Providence continues with the Star of David, then with the Pentastar, and terminates with the Octastar. The last two have only rotational symmetry. Combinatorics of Pentastar cannot be reached from Octastar by removing 33 $\triangle$. This excludes any inductive construction of the best $\bigcup$.
Moving onto $n=4$, similar computer-assisted overhaul show that $e(4)=33$ (Fig. 3). If we knew all possible arrangements of $3n$ lines in the affine plane (seq. A090338), this catalogue would encode, among other things, the Politrigon seq. (and $i(n)$, too). Such list (for $n$ lines) would also encode the solution to the Kobon $\triangle$ problem (seq. A006066). The latter asks how many non-overlapping $\triangle$ can be formed by $n$ lines (see Fig. 4 depicting one of maximal configurations for $n=10$). In its full generality, Kobon’s problem is still unsolved. It is obvious that such full register of line configurations is of no use for large $n$ do to extreme growth (of order $2^{\Theta(n\log n)}$) of their total number. In fact, the last exact value known is for $9$ lines, which amounts to $3.111.341$ simple arrangements. In our setting, this would settle, with some effort, the case $n=3$. Clearly, in all derivatives of line arrangement problem, other method should be looked for.
And it does exist in regular $\triangle\bigcup$ case. With such tools at hand, it will be possible to demonstrate the unconditional inequality $(n+1)e(n)\leq(12n-6)n$. This leads to the discovery of two “perfect" $\triangle$ $\bigcup$ showing that $e(5)=45$ and $e(8)=80$ are possible. Together with a $\triangle$ and a hexagram, this is the complete list of maximally saturated $\triangle$ $\bigcup$. In other words, the only cases where the equality is achieved (Fig. 5). (exactly the very reincarnation of two magic crosses hinted to in the Introduction). For $n\neq 1,2,5,8$, strict inequality must hold.
Theorem..
The seq. $e(n)$ starts as $3$, $12$, $22$, $33$, $45$, $56$, $67$, $80$, $91$. It also satisfies the following:
$11\leq e(n+1)-e(n)\leq 14\quad(forn\geq 3),$
$e(n)\leq 12n-18+I(10\leq n\leq 13)\quad(forn\geq 10).$
The first inequality gives situational lower bound $e(n)\geq 11n-8$ ($n\geq 10$). This can always be improved, provided we manage to find a concrete $N\geq 10$ and a configuration of NN $\triangle$ with more than 11N811N-8 edges.
The upper bound now stands in sharp contrast with a superlinear growth of the function $i(n)$. It not just bounds $e(n)$ from above, but in fact gives the exact combinatorial bound on first approximation. Without explaining this vague phrase now, here is a little trick to be played on the readers. Among all triangulations of a regular $n$-gon, choose the one whose maximal sum of number of diagonals adjacent at three consecutive vertices is minimal. Denote this minimum by $\delta(n)$ (see Fig. 6). One of the many interpretations of Catalan numbers $C_{n}=\frac{1}{n+1}\binom{2n}{n}$ is that $C_{n-2}$ gives the number of triangulations of a convex $n$-gon. Since there are n3n-3 diagonals, $\delta(n)\geq\big\lceil 6-\frac{18}{n}\big\rceil$. The following statement is therefore a result of full check ($n\leq 17$) and explicit construction ($n\geq 18$):
Proposition 1.
The sequence $\delta(n)$, $n\geq 3$, starts as $0,2,3,3,4,5,4,5,5,5,5,5$, and then stabilizes at the value $6$ (see Fig. 6).
The reader can verify that for $n\geq 2$, $\mathscr{U}(n)=12n-12-\delta(n+1)$ coincides with exact value of the seq. $e(n)$ and gives the exact upper combinatorial bound for $n\geq 10$. The relation of $\triangle$ $\bigcup$ to triangulations of $n$-gons is unexpected, and will be clarified later. But now, let us discuss what does it mean “combinatorial" vs. “geometric". [^3] Since we started from linguistics, it is apt to emphasize here that suffix dilemma ‘-ial’ vs. ‘-ic’ essentially coincides with whether it is of Latin or of Greek origin. In other words, what are pseudolines and what is stretchability.
Figure 6: Triangulations of regular polygons which minimize the maximum of number of diagonals adjacent at three consecutive vertices. This minimum is denoted by $\alpha(n)$
Pseudolines vs. Lines
Apart from A006066 and A090338, there are other seq. in the OEIS made from kindred ingredients. This means: seq., whose definition employs only well-familiar objects from Euclidean geometry, but whose calculation requisites rather delicate arrangements of these. One of such seq. was popularized by Neil Sloane and the community of OEIS during the “ranking process" invoked to clarify which seq. should be assigned a honorary label A250.000. A runner-up in this vote, now a seq. A250.001 (first proposed by Jonathan Wild), asks to find the number of ways to draw $n$ circles in the affine plane sloane ; oeis (definitions of admissible and equivalent configurations are intuitive and can be found in the references cited). In Figure 7, three (out of $173$) arrangements of $4$ circles are reproduced.
Figure 7: Three ways to draw 4 circles. Configuration $\mathbf{c}$ has $12$ boundary arcs.
Even having a ready example in front of the eyes, it takes some practice to correctly drag around centres and radii of four circles in order to achieve configurations $\mathbf{a}$ and $\mathbf{b}$. See a recent survey cutler for many more examples of such seq., associated with cutting of plane with all kinds of “exotic knives".
On the other hand, pseudoline arrangements and stretchability problem are not only close in spirit, but play a significant role in determining $e(n)$. A pseudoline is a simple connected curve in the projective plane such that its removal does not disconnect $\mathbb{R}P^{2}$. Arrangement of pseudolines is a finite set of them so that each two meet exactly once, and there are at least two such intersection points (the affine case is modified accordingly). A survey by Felsner and Goodman felsner is another enriching reading. The most transparent arrangement of projective pseudolines that cannot be realized with lines (that is, “non-stretchable") is given by a twisted Pappus configuration as in Fig. 8 (interpretation of Farey addition in terms of Pappus theorem was even featured on the cover of a recent Intelligencer schwartz ).
Figure 8: Non-stretchable arrangement of $9$ projective pseudolines
Figure 9: Non-stretchable simple arrangement of $9$ affine pseudolines
Our main concern is, however, a slightly less complicated object, simple arrangements of $n$ affine pseudolines (no multiple intersection allowed). If $n\leq 7$, any one of them can be realized with lines. Yet, the configuration in Fig. 9 (found by L. Reeves and J. Wild again) with $8$ pseudolines cannot be straightened. A best current bound on number of simple arrangements of $n$ affine pseudolines was given by Justin Dallant dallant and is $2^{0.6496n^{2}}$. It is based on Zone Theorem (that will be discussed later). Compare this bound to a one given earlier for simple arrangements of $n$ lines.
The reason for introducing pseudolines is that their arrangement problems are easier and some aspects can be tackled solely combinatorially. we will witness this. It appears that many (but not all!) results about lines are more about their configuration that straightness property itself. There are other reasons to introduce pseudolines. For example, Suvorov suvorov gives two simple arrangements of $13$ lines in the affine plane. As pseudolines, one can be transformed into the other via a continuous deformation of the plane, so that the configuration always remains simple (flexible isotopy). This, however, cannot be done if curves are always required to remain straight (rigid isotopy).
Take the $\bigcup$ of $n$ $\triangle$ labelled $1$ through $n$. Go along the border of the $\bigcup$ and write down a number of $\triangle$ to which the next edge belongs. The result of this paper amounts to constructing a seq. of maximally possible length $\mathscr{U}(n)$ which serves as a model of such actual seq. The second step would be to show that this model is realisable in pseudoline geometry. Possibly, this is not too hard to show. The question whether a construction can further be straightened is of a very different sort. Stretchability problem is known to be NP-hard.
Lower Bound
It is straightforward to contrive $n\geq 1$ $\triangle$ whose $\bigcup$ is a simple polygon with 9n69n-6 edges. One only needs to choose $3$ points on $3$ different sides of the existing polygon, slightly protrude the points outwards, and construct a new $\triangle$ (Fig. 10A). The only two concerns are that no vertex of an initial polygon lie inside the newly constructed $\triangle$, and that the new polygon is simple. If one vertex of the initial polygon with reflex ($180^{\circ}$) angle lies inside a $\triangle$, we get an increase of edge count by $10$ (Fig. 10B). If two non-consecutive vertices with reflex angles are inside, the increase is by $11$ (Fig. 10C).
Figure 10: Adding one $\triangle$ changes an edge count.
With this idea in mind, let us take, as a start-up, three brown $\triangle$ as shown in Fig. 11. Their $\bigcup$ has $22$ edges. Consider a blue segment and two green circles as shown. The latter are inscribed into $\angle PXA$ and $\angle SYB$, each passing through points $A$ and $B$, respectively. Next, choose several points (labelled $1,2,\ldots,m=n-3$, left to right) strictly inside the blue segment. From every point $p$, draw “inner" tangents to both circles. Let the intersection with $PQ$ and $SR$ be, respectively, $p^{\prime}$ and $p^{\prime\prime}$. Fig. 11 shows example with $m=2$.
Figure 11: An inductive construction showing that $e(n)\geq 11n-11$ for $n\geq 4$.
We now miniaturelly protrude all points on $PQ$, SRSR and the blue segment to the exterior. Each of $m$ $\triangle$ adds $11$ edges. This gives the total count of 11n1111n-11.
In fact, nothing is specific about this start-up configuration. The construction works each time there exists an edge which is not adjacent to any vertex of an original $\triangle$ (so, $n\geq 3$). Take, for example, the green edge in Fig. 12 (the one with a control vertex numbered “47"). Its endpoints belong to a cyan and grey $\triangle$, respectively. Take a point $x$ inside this edge. We will construct a $\triangle$ whose vertex is $x$, only slightly protruded outwards. It will lie between black and brown vertices $a$ and $b$, respectively. Let left and right edges at these vertices intersect at point $p$ and $q$ (shown also on the same Fig. 12). Draw line $xp$ and $xq$ and find their intersections $y,z$ with the corresponding brown (shown in bold) and black segments on the boundary of the union. Slightly inflated $\triangle$ xyzxyz adds +11+11 in edge count. This mimics precisely the construction in Fig. 11. Also, vertices of triangles preserve a cyclic order.
Figure 12: $\bigcup$ of 88 $\triangle$ with $79$ edges, whose seq. is given by (1) (ignore $\triangle xyz$ for now). All vertices of 88 $\triangle$ are on the unit circle; few control vertices of a $79$-gon are labelled. $\triangle xyz$ is used in the inductive prove to show that $e(8+k)\geq 79+11k$.
Note that such lower bounds are situational. To see this, consider an Octastar in Fig. 5. Removing any 33 $\triangle$ give a collection of 55 $\triangle$ with at most $44$ edges in the $\bigcup$. In other words, we cannot expect to achieve maximum number of edges for $(n+1)$ $\triangle$ by finding the best addition to the existing best $n$ $\triangle$ configuration. Each $n$ should be treated individually, what we will do in the section ‘Maximal Sequences’.
Upper Bound
Take a regular $\bigcup$ of $n$ $\triangle$ with $M$ edges. For example, consider 88 $\triangle$ in Fig. 12. We start with a blue $\triangle$ (labelled $1$) and an edge just before the vertex with control label “10". All $\triangle$ are numbered $1$ through $8$. So, for example, black $\triangle$ gets number $4$. Now, go along the boundary in a positive direction and record what edges occur. This produces the cyclic seq. of length $79$:
$\underline{\underline{11}{\color[rgb]{1,0,0}3}\overline{2233{\color[rgb]{1,0,0}216}4455{\color[rgb]{1,0,0}4}66{\color[rgb]{1,0,0}418}7788{\color[rgb]{1,0,0}7}11}}\overline{{\color[rgb]{1,0,0}74}22}{\color[rgb]{1,0,0}4}3344{\color[rgb]{1,0,0}3276}$
$5566{\color[rgb]{1,0,0}5}77{\color[rgb]{1,0,0}521}88\underline{\underline{11{\color[rgb]{1,0,0}8}22{\color[rgb]{1,0,0}85}33{\color[rgb]{1,0,0}5}4455{\color[rgb]{1,0,0}4387}6677{\color[rgb]{1,0,0}6}88{\color[rgb]{1,0,0}63}}}.$ (1)
Take a fragment of the boundary contained between two vertices of the same blue $\triangle$. Say, the simply-underlined one in (1). Write down numbers of $\triangle$ whose RIGHT edges are encountered along the way. That is: ignore a number if its double has not yet occurred; otherwise, include it; from each doubled number, only one copy is included. This way the seq. $\mathcal{L}=123214546417871$ is obtained. What properties are intrinsic to it? First, no two adjacent numbers are equal. Second, though some of them may (and do) appear multiple times, such seq. do not contain a subseq. of the form abababab for any distinct pair $a,b\in\{1,2,\ldots,n\}$. This is a crucial property of regular $\triangle$ $\bigcup$. In the section ‘Lower envelopes’ we will see that as long as such seq. are allowed to have subseq. of the form abababab, their length can grow superlinearly.
How to prove the stated property? Since we go along the common boundary counter clock-wise, each leg has a direction. Drawing a full line through its endpoints, we can talk about right and left side of it. The same applies to any chosen edge. All the remaining edges split into lying to the right, to the left of it, or collinear. Also, for any two points on the same leg, we can say which one is behind and which one is ahead the other. Now, fix any number in the produced seq. of right edges. All the successive edges are either collinear (those with the same number), or lie to the right of it. Suppose, a subseq. abaaba belongs to such a seq. Let legs $a$ and $b$ intersect at a point $P$. In the process of recording right edges, we pass from leg $a$ to leg $b$, and then to leg $a$ again. Stop here. The point $P$ now lies behind. It is no longer possible to get beck to leg $b$, since all points of leg $b$ ahead of $P$ lie to the left of leg $a$.
Hence, $\mathcal{L}$ and all such seq. are exactly what is known to be a $(2,n)$-Davenport-Schinzel (or DS) seq. by definition sharir . The maximal length of $(2,n)$-DS is known to be 2n12n-1.
The proof of this fact is done by induction. Suppose, it is true for n1n-1. Consider a $(2,n)$-DS seq. $\mathcal{L}$. If not all symbols $1$ through $n$ appear, it is then a $(2,m)$-DS seq. for some $m<n$. Otherwise, reading left-to-right, assume that symbols appear in their natural order: $1,2,\ldots,n$ (one can always rename if needed). We claim that symbol $n$ appears in $\mathcal{L}$ only once. Assuming the opposite, let $\mathcal{L}$ contain a subseq. npnnpn for some $p\leq n-1$. Due to natural order convention, $\mathcal{L}$ must also contain a subseq. pnpnpnpn – a contradiction. Removing $n$ and concatenating two symbols which surround it in case they are equal, we get a (2,n1)(2,n-1)-DS seq. of a maximal length 2n32n-3. Thus, $\mathcal{L}$ has a length at most 2n3+22n-3+2. Note also that in case $(2,n)$-DS seq. is maximal, the last symbol to appear must always be surrounded by two identical symbols.
Let us return to our edge count. Take the same underlined fragment, but start from the other end. That is, write down numbers of $\triangle$ whose LEFT edge contribute to the boundary. This way, seq. $18781654613231$ is obtained. It is again a $(2,n)$-DS seq. Consequently, the total number of edges between two blue vertices is $\leq 4n-2$, with a stipulation that all blue edges were counted twice. We perform the same calculation for all three blue vertex pairs. Thus, if $\ell_{i}$ is the number of edges the $i$-th $\triangle$ contributes to the boundary of $M$-gon, then
$2\ell_{1}+\ell_{2}+\cdots+\ell_{n}\leq 12n-6.$ (2)
Summing over all $i$ yields $(n+1)M\leq n(12n-6)$. Thus, $M\leq 12n-18+\frac{18}{n+1}$. This looks great, with a caveat that in case $n=7$, the bound stands at $68$, while $67$ is the best we have yet managed to achieve $(45+11+11)$.
The inequality $e(n)<12n$ also implies the following. Take a regular $\bigcup$ of $n$ $\triangle$ with $e(n)$ edges. Choose $\triangle$ which contributes least to the common boundary. Its toll is thus $\leq 11$ edges. Remove this $\triangle$ from the $\bigcup$. Each edge not adjacent to its vertex reduces an edge count by $1$, while two edges at any vertex reduce this by $3$. Therefore, $e(n-1)\geq e(n)-14$.
The idea to separate left and right edges between two same-colored vertices was inspired by the proof of the aforementioned Zone Theorem given by Edelsbrunner at al. in edelsbrunner . In the latter paper though, ‘right’ and ‘left’ correspond to two directions along the same ray. Among several existing proofs of the Zone Theorem, it is the one which essentially employs $(2,n)$-DS sequences.
What this result claims? Consider $\mathcal{L}$, a simple arrangement of $n$ lines in the affine plane. The set $\mathbb{R}^{2}\setminus\mathcal{L}$ split into connected components called faces, which are polygons (in case face is bounded) or faces with two infinite edges. The size of a face is the number of its edges. Given a line $\ell$. If $\mathcal{L}\cup\{\ell\}$ is a simple arrangement, the zone of $\ell$ is a collection of all faces it crosses. The complexity of the zone of $\ell$ is the sum of sizes of all faces in the zone. This notion was first introduced by Chazelle at al. in zone , where an upper bound $6n$ for complexity was given. DS seq. approach yields 6n46n-4 (see edelsbrunner , p. 334, also sharir , p. 125). In horizon , Bern at al. strengthened this to $\lfloor 5.5n\rfloor-1$. Rom Pinchasi pinchasi later improved it to $\lfloor 5.5n\rfloor-3$ pinchasi . A construction in Fig. 13 (taken from horizon , just slightly enhanced) shows that the latter bound is almost optimal. Note the similarity with construction in Fig. 11.
Figure 13: An arrangement of $11$ lines; a complexity of the zone of blue line (colored in green) is $54=\lfloor 5.5\cdot 11\rfloor-6$; a size of each face in the zone is marked.
Maximal Sequences
The final insight comes from the observation that, depending on $n$, a certain amount of all $6n$ of the obtained seq. must be non-maximal. In terms of the upper bound, this only settles the case $n=7$, and reduces it by $1$ for $15\leq n\leq 18$. This seems a feeble contribution. However, such deeper analysis also gives an algorithm to construct a seq. of length $\mathscr{U}(n)$ which does not violate combinatorics of pseudolines.
For the starters, let us return to the example $\mathcal{L}_{1}=123214546417871$. Its length is $2\cdot 8-1$. So, it is maximal ($\mathbf{M}$ for short). Now, shift our attention to the cyan triangle. Or, better, the overlined fragment in (1). The produced seq. is also M. We argue that this extension to $\mathcal{L}_{1}$ can be reproduced from $\mathcal{L}_{1}$ itself. Indeed, starting from $\mathcal{L}_{1}$, remove all (but the last) $1$’s and place $2$ in the end. This results in $232454647871\ldots 2$. For this to be $\mathbf{M}$, dots should be replaced by two numbers: $\mathcal{L}_{2}=232454647871ab2$. We claim that $a,b$ can be uniquely reconstructed from a prefix $232454647871$ and the terminal number $2$. The argument behind this will now certainly look familiar. Let us methodically remove the last (reading left to right) symbol to appear, and merge its two neighbours. Such merge must always occur:
$2324546478\underline{71a}b2$ $\mathop{\longrightarrow}^{a=7}\,\,23245464\underline{787}b2\rightarrow$
$2324546\underline{47b}2$ $\mathop{\longrightarrow}^{b=4}\,\,232454642.$
This gives the only possible reconstruction: $\mathcal{L}_{2}=232454647871742$. The overlined fragment in (1) confirms it. In general, a completely identical argument proves the following statement.
Lemma.
Given a prefix $\mathcal{P}=1\ldots 2\ldots 3\ldots n$ and the terminal number $1$. If $\mathcal{P}1$ is a $(2,n)$-DS seq., there exists the unique $\mathcal{Q}$ such that $\mathcal{P}\mathcal{Q}1=\mathscr{S}(Q)$ is maximal.
This allows to iterate the procedure (reconstructed symbols are printed in red):
$\scriptstyle 123214546417871\rightarrow 232454647871{\color[rgb]{1,0,0}74}2\rightarrow 3454647871742{\color[rgb]{1,0,0}4}3\rightarrow$
$\scriptstyle 454647871742434\rightarrow 5678717234{\color[rgb]{1,0,0}3276}5\rightarrow 678717234327656\rightarrow$
$\scriptstyle 7871723432756{\color[rgb]{1,0,0}5}7\rightarrow 81234325657{\color[rgb]{1,0,0}521}8\rightarrow$
$\scriptstyle 123432565752181\rightarrow 2343256575281{\color[rgb]{1,0,0}8}2.$ (3)
This exactly (!) corresponds to actual seq. extracted from (1). So far, all of them are $\mathbf{M}$. We also notice that $\mathscr{S}^{9}(\mathcal{L})$ is obtained from $\mathcal{L}_{1}$ by increasing all terms by 1. As we will soon see, this is a general rule.
Let us now continue not with iterations of $\mathscr{S}$, but with actual sequences extracted from (1). For a while, $\mathbf{M}$ seq. continue. However, at a step $17$ (double underline), a nonmaximal sequence $12345436768{\color[rgb]{1,0,0}63}1$ (length $14$) occurs. Then it is followed with a $\mathbf{M}$ $2345436768631{\color[rgb]{1,0,0}3}2$. Next $2$ seq. are also $\mathbf{M}$. Then comes another shorther $56768612321645$ (length $14$), after that comes $\mathbf{M}$ $6768612321645{\color[rgb]{1,0,0}4}6$. Finally, after two more steps, everything closes up at the initial $1232145464178{\color[rgb]{1,0,0}7}1$. The conclusion: process almost always runs smoothly (and the next seq. is deterministic), only during $24$ steps two glitches occur. The total number of edges in the $\bigcup$ equals the number of red symbols in this sequence of sequences (each of length 15 or 14), plus $6\cdot 8$.
If the reader has already decided that this is too confusing to follow, this is true. Seeing such an arbitrary-looking collection of numbers in (3), it is impossible to fathom what is actually happening. In fact, with a help a visualisation of maximal $(2,n)$-sequences in terms of triangulations of regular $(n+1)$-gons (first discovered by D. Roselle roselle ; see also sharir , p.250-232), we will see that… almost nothing is happening!
To witness this, consider, as an example, a maximal $(2,9)$-DS sequence $12343536768632191$. First, divide it into $9$ decreasing portions as follows: | 1| 2| 3| 43| 53| 6| 76| 86321| 91||\,1|\,2\,|\,3\,|\,43\,|\,53\,|\,6\,|\,76\,|\,86321\,|\,91\,|. Given regular $(n+1)$-gon. For each portion starting from $i$, connect vertices with these particular labels to the vertex $i+1$. This way Fig. 14 is obtained.
Figure 14: A triangulation of a regular $10$-gon which corresponds to a maximal $(2,9)$-DS sequence $12343536768632191$.
It is straightforward to see that this not only gives the visualisation, but in fact a bijection between all triangulations of $(n+1)$-gon and essentially distinct maximal $(2,n)$-DS sequences (those which cannot be made equal after redefining or reordering initial set of symbols).
In our case, this triangulation convention is used for the natural order: $1,2,\ldots,n$. If the order is $i,i+1,\ldots i-1$, we adopt the following rule. First, rotate symbols cyclically, so that the order is $1,2,\ldots,n$. So, for $\mathscr{S}(\mathcal{L}_{1})=232454647871742$ one gets $121343536768631$. Now, use the standard convention for visualisation, only the distinguished yellow vertex should be moved to position $i$.
Let as apply this rule for visualising actual seq. extracted from (1). The result is presented as Fig. 15. It is surprising and relieving at the same time: nothing changes going from seq. 1 to seq. 2, except that the distinguished vertex moved one position. In light of such triangulation correspondence, Lemma become a tautology. Given a triangulation of a regular $n$-gon. Then it tells that if one removes all edges ending at one particular vertex, then there is only one way to complete the triangulation by adding edges starting from this particular vertex.
Such pattern continues up to seq. 8. Going to seq. 9 (it starts again at 1), the whole triangulation rotates one turn in the positive direction. This becomes clear if given some thought.
Now, what happens when a glitch (a shorter seq.) occurs? How it is succeeded by a $\mathbf{M}$ seq.? The same Fig. 15 (in the traditions of Intelligencer) might be considered as a proof without words of the following fact: at step $16\rightarrow 17$ (distinguished is vertex 8), one edge ending with $8$ is removed (at the same time, since the vertex changes from $8$ to $1$, a one-step rotation in positive direction occurs). At step $17\rightarrow 18$ (vertex 1) and edge with endpoint $1$ is added. After two more moves, another such incident takes place: one edge is removed, a new edge is reinstated during the next move.
We come to understand that the whole process runs as follows. Start from a distinguished vertex in a triangulation of a regular $(n+1)$-gon. Perform $3n$ moves. An $i$th move amounts to removing diagonals at this vertex; adding other diagonals (at the same vertex); counting their number $t_{i}$; switching to the next neighbour. After the whole operation, vertex n1n-1 is reached. The goal is to obtain the same triangulation, only rotated $3$ turns in the negative direction. The total combinatorial length of such $\triangle$ $\bigcup$ is then $\sum_{i=1}^{3n}t_{i}+6n$.
Figure 15: Triangulations corresponding to actual seq. in of RIGHT edges in Fig. 12; 1-9 also correspond exactly to seq. in (3)
If no deletions are necessary, it means that the triangulation coincides with itself if rotated by 3-3 moves. So, $\alpha(n)$ equals $6$ times the total number of edges in the triangulation, that is, 6(n+1+n2)=12n66(n+1+n-2)=12n-6, minus the sum of number of edges at three consecutive vertices. For $n+1=6$ and $9$ we see that this sum is equal to $9$ and $10$, respectively, independently of the start vertex. As expected, this gives $12\cdot 5-6-9=45$ and $12\cdot 8-6-10=80$ edges of the $\bigcup$.
Lower Envelopes
This section is very sketchy, mostly referring to the monograph of Sharir and Agarwal sharir . Take $7$ linear univariate functions defined on the real line. Label them $1$ through $7$. Let $f$ be defined as the minimum of these. If, while taking the minimal value, we jump from one line to the other, it is obviously no longer possible to go back to the abandoned one. Therefore, the support sequence (telling which function provides the minimum as one progresses along the $x$-axis) for $f$ consist of at most $7$ symbols.
However, if each linear function is defined on its own interval so that their $\bigcup$ is also an interval, the support sequence can be longer. This is clear. What is astonishing that it can be much, much, longer! As an example, consider $7$ segments as in Fig. 16.
Figure 16: The support for lower envelope of 77 segments with complexity $25$
The support seq. $1213145416164653737572767$ has length $25$. What properties does such seq. have? Similarly as with $\mathcal{L}$, no two adjacent symbols are equal. At the same time, we immediately notice its subseq. like $1313$, $1414$, $3535$, etc. This is a major distinction. We leave it as an exercise for th reader to prove that though subseq. abababab may be present, subseq. of the for ababaababa are not to be found. This is exactly the definition of a $(3,n)$-Davenport-Schintzel seq.
The foundational part of sharir (p. 12-42) is devoted to investigating $(3,n)$-DS seq. The story how their true complexity unravelled is inspiring. The final answer – the maximal length of such seq. is $2n\alpha(n)+O(n)$, a result by Seth Pettie pettie – tells little about the story. In the preface to the book is is recalled how Hart and Sharir went onto the discovery that superlinear grown is possible.
Here is another major achievement. Consider a $(3,n)$-DS, now of maximal possible length $27$ (see seq. A002004 in OEIS): $121314151616543272737475767$. It can be realized as a lower envelope of $7$ pseudosegments (connected parts of $7$ pseudolines). Can it be realized as a lower envelope of $7$ segments? In other words, can pseudosegment construction be stretched? I do not know. What is certain that there is little hope (and most likely, untrue; see sharir , p. 112) that a certain maximal $(3,n)$-DS seq. for large $n$ can serve as an envelope of $n$ segments. However, Wiernik and Sharir proved that there exist $(3,n)$-DS seq. whose length asymptotically differs by a factor of at most $2$ from the length of longest such seq. and which does serve as lower envelope of $n$ segments. Speaking philosophically, this shows that the Ackermann function, first discovered in logic and computability analysis, is intricately present in nature, if we assume that straight segments are. Along with Turing machines with record-halting time strauss , aperiodic hat monotile, tiles with high Heesch number, this is something what carries mathematics well outside the domain of natural sciences.
To finish the construction, consider Fig. 16 once again. If we imagine each segment completed to a triangle by two nearly vertical legs, we clearly can get a simple polygon. Thus, finding $i(n)$ rests on other fundamental problems about unions. The following problem might serve a good training field:
Problem.
Find $n$ for which $i(n)>e(n)$.
Outro
During the Journées Arithmétiques conference held in Vilnius, Lithuania in 2011, Artūras Dubickas gave a talk. Andrzej Schinzel, then in the audience, noted the question posed by the speaker. Later he wrote a note containing the solution (Fig. 17). That is how an autograph of a famous mathematician landed in a country-side of Northern Lithuania. Though it is a bit of a stretch, but numbers $n=1,2,5,8$ which give rise to perfect trangle unions are also in the Fibonacci sequence.
Figure 17: A handwritten note made by Andrzej Schinzel in Vilnius, 2011